# Enabling Auditory Large Language Models for Automatic Speech Quality Evaluation

### v1

@@ Abstract
Speech quality assessment typically requires evaluating audio from multiple aspects, such as mean opinion score (MOS) and speaker similarity (SIM) etc., which can be challenging to cover using one small model designed for a single task. In this paper, we propose leveraging recently introduced auditory large language models (LLMs) for automatic speech quality assessment. By employing task-specific prompts, auditory LLMs are finetuned to predict MOS, SIM and A/B testing results, which are commonly used for evaluating text-to-speech systems. Additionally, the finetuned auditory LLM is able to generate natural language descriptions assessing aspects like noisiness, distortion, discontinuity, and overall quality, providing more interpretable outputs. Extensive experiments have been performed on the NISQA, BVCC, SOMOS and VoxSim speech quality datasets, using open-source auditory LLMs such as SALMONN, Qwen-Audio, and Qwen2-Audio. For the natural language descriptions task, a commercial model Google Gemini 1.5 Pro is also evaluated. The results demonstrate that auditory LLMs achieve competitive performance compared to state-of-the-art task-specific small models in predicting MOS and SIM, while also delivering promising results in A/B testing and natural language descriptions. Our data processing scripts and finetuned model checkpoints will be released upon acceptance.
@@ II-B Automatic Speech Quality Evaluation
Recent advancements in automatic speech quality evaluation have primarily focused on predicting MOS, which is often obtained by averaging discrete opinion scores assigned by dozens of independent human listeners. AutoMOS [1] employs a deep recurrent neural network to predict MOS, achieving correlations closely aligned with human evaluators. MOSNet [2] explores different model architectures, identifying CNN+BLSTM as the top performer. SSL-MOS [4] further enhances accuracy by utilizing pre-trained speech SSL models. UTMOS [3], based on ensemble learning of strong and weak learners, achieved the highest scores on several metrics in the main track of the VoiceMOS Challenge 2022 [28]. Moreover, the recently developed speech language models have been applied to provide alternative perspectives on synthesized speech evaluation, showing promising correlations with subjective metrics [29, 30].
@@ III-A Auditory LLMs for Audio Perception and Understanding
SALMONN [8] has a dual-encoder architecture, with Whisper-Large-v2 [35] encoder as speech encoder and BEATs encoder [36] as non-speech audio encoder. This design enables SALMONN to effectively interpret a wide range of general audio inputs. The outputs from both encoders are concatenated and fed into a connection module, a window-level Q-Former, which shares the same architecture as the vanilla image Q-Former [37] but operates at the window level. Instead of processing the entire audio at once, the window-level Q-Former generates one textual token for approximately every 0.33 seconds, resulting in 88 tokens for a 30-second audio. The LLM backbone is the Vicuna LLM [38] adapted by LoRA, which generates responses with audio textual tokens and text prompt tokens presenting.
@@
TABLE I: MOS prediction results on test split of NISQA, BVCC and SOMOS datasets. For NISQA, the average results of the three public available test sets (NISQA_TEST_LIVETALK, NISQA_TEST_FOR and NISQA_TEST_P501) are reported. No system-level results for NISQA because there is no system label in NISQA dataset.
@@ III-B Speech Quality Evaluation with Auditory LLM
  • MOS prediction: Auditory LLMs predict the MOS score on a scale of 1 to 5, with precision to one decimal place, for a given audio input. To improve performance across different datasets, dataset-specific prompts are designed. These prompts distinguish between datasets by incorporating the phrase “according to Dataset standards”, allowing the model to adjust its predictions based on the particular characteristics of each dataset.
@@ IV-B Models
Four auditory LLMs are used in this paper, including SALMONN (vic1.5), SALMONN (vic1.0) [8], Qwen-Audio [9] and Qwen2-Audio-7B [11]. The LLM backbones used by the SALMONN models are denoted in parentheses, where vic1.5 denotes Vicuna-v1.5-7B and vic1.0 denotes Vicuna-v1.0-13B. These auditory LLMs are finetuned using LoRA [34] with rank of 8 and scale of 32, across all four tasks outlined in Sec. IV-A, following the official scripts released [^1] For SALMONN: https://github.com/bytedance/SALMONN; For Qwen-Audio: https://github.com/modelscope/ms-swift/issues/1653. The model is trained for 10 epochs, and the one that performs best on the validation set is selected for testing.

### v2

@@ Abstract
Speech quality assessment typically requires evaluating audio from multiple aspects, such as mean opinion score (MOS) and speaker similarity (SIM) etc., which can be challenging to cover using one small model designed for a single task. In this paper, we propose leveraging recently introduced auditory large language models (LLMs) for automatic speech quality assessment. By employing task-specific prompts, auditory LLMs are finetuned to predict MOS, SIM and A/B testing results, which are commonly used for evaluating text-to-speech systems. Additionally, the finetuned auditory LLM is able to generate natural language descriptions assessing aspects like noisiness, distortion, discontinuity, and overall quality, providing more interpretable outputs. Extensive experiments have been performed on the NISQA, BVCC, SOMOS and VoxSim speech quality datasets, using open-source auditory LLMs such as SALMONN, Qwen-Audio, and Qwen2-Audio. For the natural language descriptions task, a commercial model Google Gemini 1.5 Pro is also evaluated. The results demonstrate that auditory LLMs achieve competitive performance compared to state-of-the-art task-specific small models in predicting MOS and SIM, while also delivering promising results in A/B testing and natural language descriptions. Our data processing scripts and finetuned model checkpoints can be found at https://github.com/bytedance/SALMONN.
@@ II-B Automatic Speech Quality Evaluation
Recent advancements in automatic speech quality evaluation have primarily focused on predicting MOS, which is often obtained by averaging discrete opinion scores assigned by dozens of independent human listeners. AutoMOS [1] employs a deep recurrent neural network to predict MOS, achieving correlations closely aligned with human evaluators. MOSNet [2] explores different model architectures, identifying CNN+BLSTM as the top performer. SSL-MOS [4] further enhances accuracy by utilizing pre-trained speech SSL models. UTMOS [3], based on ensemble learning of strong and weak learners, achieved the highest scores on several metrics in the main track of the VoiceMOS Challenge 2022 [28]. Moreover, the recently developed speech language models have been applied to provide alternative perspectives on synthesized speech evaluation, showing promising correlations with subjective metrics [29, 30, 31, 32].
@@ III-A Auditory LLMs for Audio Perception and Understanding
SALMONN [8] This design enables SALMONN to effectively interpret a wide range of general audio inputs. The outputs from both encoders are concatenated and fed into a connection module, a window-level Q-Former, which shares the same architecture as the vanilla image Q-Former [37] but operates at the window level. Instead of processing the entire audio at once, the window-level Q-Former generates one textual token for approximately every 0.33 seconds, resulting in 88 tokens for a 30-second audio. The LLM backbone is the Vicuna LLM [38] adapted by LoRA, which generates responses with audio textual tokens and text prompt tokens presenting.
@@
TABLE I: MOS prediction results on test split of NISQA, BVCC and SOMOS datasets. For NISQA, the average results of the three public available test sets (NISQA_TEST_LIVETALK, NISQA_TEST_FOR and NISQA_TEST_P501) are reported. No system-level results for NISQA because there is no system label in the NISQA dataset.
@@ III-B Speech Quality Evaluation with Auditory LLM
  • MOS prediction: Auditory LLMs predict the MOS score on a scale of 1 to 5, for a given audio input. To improve performance across different datasets, dataset-specific prompts are designed. These prompts distinguish between datasets by incorporating the phrase “according to Dataset standards”, allowing the model to adjust its predictions based on the particular characteristics of each dataset.
@@ IV-B Models
Four auditory LLMs are used in this paper, including SALMONN (vic1.5), SALMONN (vic1.0) [8], Qwen-Audio [9] and Qwen2-Audio-7B [11]. The LLM backbones used by the SALMONN models are denoted in parentheses, where vic1.5 denotes Vicuna-v1.5-7B and vic1.0 denotes Vicuna-v1.0-13B. The connectors and LoRA on LLM backbones are finetuned, across all four tasks outlined in Sec. IV-A, following the official scripts released [^1] For SALMONN: https://github.com/bytedance/SALMONN; For Qwen-Audio: https://github.com/modelscope/ms-swift/issues/1653. The LoRA rank is 8 and the scale is 32. The model is trained for 10 epochs, and the one that performs best on the validation set is selected for testing.